One kg of fluid enters the steady flow apparatus at a pressure of P1, velocity 16m/s and

specific volume 0.4 m3/kg. The Inlet is 30 m above the ground level. The fluid leaves the

apparatus at Pressure of P2, velocity of 275 m/s and specific volume 0.6 m3/kg. The outlet is

at the ground level. The total heat loss between the inlet and outlet is 10 KJ/kg of fluid. If 140

KJ/kg of work is done by the system, find the change in specific internal energy and indicate

whether this is a increase or decrease.

P1 = 6 Bar − {

√X5

849 } Pa ; P2 = 1 Bar + {+√99 − X

4

}Pa

x= 51


1
Expert's answer
2021-12-24T13:43:03-0500

The steady flow energy equation:

Q W = mΔH + 1/2m(C22 − C12) + mg(Z2 Z1)

ΔH = (U2 + p2"\upsilon"2) − (U1 + p1"\upsilon"1) = ΔU + p2"\upsilon"2p1"\upsilon"1

"C" - velocity, "p" - pressure, "v" - specific volume, "Z" - height, "U" - internal energy.

10 x 103 J/kg - (-140 x 103 J/kg) = 1 kg x ΔH + 1/2 x 1 kg x ((16 m/s)2 − (275 m/s)2) + 1 kg x 9.81 m/s2 x (30 m − 0 m)

150 x 103 J/kg = ∆H - 37 390 J/kg

H = 187 390 J/kg

p2 = 1 Bar + {+√99 − X4}Pa = 1 Bar + {+√99 − 4 x 51} = 100 000 Pa - 194 Pa = 99 806 Pa

p1 = 6 Bar − {√X5849} Pa = 600 000 Pa - {√51 x 5849} Pa = 600 000 Pa - 546 Pa = 599 453 Pa

187 390 J/kg = ∆U + 99 806 Pa x 0.6 m3/kg - 599 453 Pa x 0.4 m3/kg

U = 367 288 J/kg = 367.3 kJ/kg (positive sign indicates that specific internal energy increased)